OpenStudy (anonymous):

A ball is tossed from an upper-story window of a building. The ball is given an initial velocity of 8.20 m/s at an angle of 17.0° below the horizontal. It strikes the ground 6.00 s later. (a) How far horizontally from the base of the building does the ball strike the ground? m (b) Find the height from which the ball was thrown. m (c) How long does it take the ball to reach a point 10.0 m below the level of launching?

OpenStudy (jfraser):

break the velocity vector into horizontal and vertical components solve each one independently using the velocities and time

OpenStudy (irishboy123):

a) horizontal velocity = 8.2cos(17) = 7.84m/s horizontal distance travelled in 6 secs = 6 * 7.84 = 47m b) vertical velocity = 8.2sin(17) = 2.4m/s x = ut + 1/2a(t*t) = 2.4*6 + 1/2(9.8)(6*6) = 190.8m c) you could use x = ut + 1/2at^t BUT to avoid more painful quadratic solve it this way: vertical velocity after it has dropped 10 m given by: v^2 = u^2 + 2as v^2 = 2.4*2.4 + 2(9.8)(10) = 201.8 v = 14.2m/s then from v = u + at t = (v-u)/a = (14.2-2.4)/9.8 = 1.2s.